# Operator Deep Smoothing for Implied Volatility

### v2

@@ 5 Discussion
We provide a novel method for implied volatility smoothing, resulting from an application of our general operator deep smoothing approach for discretization-invariant data interpolation. The approach leverages a graph neural operator to directly map given data – consistently across size and spatial arrangement – to smoothed surfaces, transcending classical parametric smoothing techniques. In the example of volatility smoothing, benefits include a massively simplified online calibration process.
@@
By moving the application of neural networks from the instance-by-instance level (Ackerer et al.,, 2020) to the “operator level”, we leverage the information contained in the entire training dataset for the smoothing of every single surface. In other words, our method “unlocks” large historical options datasets for volatility smoothing. We argue that our substantial outperformance against Ackerer et al., (2020) in the “Extrapolation-Test”-setting of the benchmark detailed in Table 5 of Appendix C.5 owes to this circumstance.
@@
Figure 3 makes the compression qualities of the operator deep smoothing approach apparent: We compute the entire historical timeseries using a single GNO instance, with around 100 thousand parameters. Evaluating the SVI benchmark, on the other hand, requires 61454 model instances (one per slice), or a total of 307270 parameters. A comparison with Ackerer et al., (2020), which for each smoothed surface trains a new neural network of around 5085 parameters, [^11] Computed as the sum of $120=3\times 40$ parameters for the input layer, three times $1640=41\times 40$ parameters for the hidden layers, $41$ parameters for the output layers, plus $4$ additional parameters of the SSVI prior and a scaling parameter. is striking: Smoothing of the CBOE dataset 2012–2021 at its 20-minute interval frequency would require more than 200 million parameters (more with rising frequency). At the same time, we expect our GNO to perform accurate smoothing over the entire training period and beyond (with regular finetuning), and our model instance remains fixed, even when moving to higher-frequency data.
@@
Compared to ad-hoc volatility parametrizations like SVI, the operator deep smoothing approach loses interpretability of parameters, which for some practitioners may be a stringent requirement. This disadvantage is generally shared by neural network based engineering solutions. Moreover, in some situations dimensionality reduction (even without interpretability of parameters) may be a desirable additional feature that is not directly achieved by our operator deep smoothing approach. Combining the VAE method (Bergeron et al.,, 2021) with our operator deep smoothing approach could lead to further promising potential applications of neural operators. Huang et al., (2024) introduces neural mappings, which generalize neural operators to mixed infinite-/finite-dimensionality for input or output spaces. This motivates a discretization-invariant GNO-based encoder, fit to handle raw incoming market data, and a classical decoder to extend the operator deep smoothing approach to a VAE-like architecture.
@@ Acknowledgements
Access to data was made possible through the Clearify Project funding by an Imperial College SME engagement grant jointly with Zeliade Systems, as well as the MSc in Mathematics and Finance, Imperial College London. AJ gratefully acknowledges financial support from the EPSRC grants EP/W032643/1 and EP/T032146/1. Computational resources and support were provided by the Imperial College Research Computing Service (Harvey,, 2017). For the purpose of open access, the author(s) has applied a Creative Commons Attribution (CC BY) licence (where permitted by UKRI, ‘Open Government Licence’ or ‘Creative Commons Attribution No-derivatives (CC BY-ND) licence’ may be stated instead) to any Author Accepted Manuscript version arising.
@@ Appendix A Neural Operators
We give a review of the kernel integral transform neural operator framework of Kovachki et al., (2023), and expand in more detail on its graph neural operator.
@@
Let $\mathcal{A}$ and $\mathcal{U}$ be input and output space of an operator learning problem, as introduced in Section 3.1. Then, $\mathcal{A}$ and $\mathcal{U}$ are Banach spaces of functions $D\to\mathbb{R}^{c_{\text{in}}}$ and $D\to\mathbb{R}^{c_{\text{out}}}$, respectively, where $D$ is a (bounded) domain in $\mathbb{R}^{d}$. The mathematical analysis of neural operators in Kovachki et al., (2023) summarized hereafter, as well as the definition of universality and discretization-invariance in Section 3.1, make use of the following terms and notations.
In the context of Kovachki et al., (2023), a domain is a bounded and connected open set that is topologically regular (in the sense that it is the interior of its closure). A domain $D$ is Lipschitz if its boundary locally is the graph of a Lipschitz continuous function defined on an open ball of $\mathbb{R}^{d-1}$. An open ball – for any metric space $\mathcal{X}=(\mathcal{X},d)$ – is the set
@@ A.1 Kernel Integral Neural Operators
Theorem A.1 (Universal Approximation; Theorem 11 of Kovachki et al., (2023)).
@@
Kovachki et al., (2023) establishes the following.
Theorem A.2 (Discretization Invariance; Theorem 8 of Kovachki et al., (2023)).
Let $F^{\theta}\colon\mathcal{A}\to\mathcal{U}$ be a kernel integral neural operator, where $\mathcal{A}$ and $\mathcal{U}$ both continuously embed into $C(\overline{D})$. Then, the implementation of $F^{\theta}$ based on (12) is discretization-invariant as defined in Section 3.1.
(12) suggests the straightforward (quasi) Monte-Carlo inspired implementation
@@ A.2 Graph Neural Operators
The curse of dimensionality makes the direct implementation (13) prohibitively expensive in practice. Instead, Anandkumar et al., (2020) introduces graph neural operators (or, GNOs, for short) which replace the kernel integral operation at the heart of the framework by a sum approximation and organizes the constituent terms using a directed graph structure: The discretization $\pi=\{x_{1},\dots,x_{m}\}$ of the input data $\mathbf{h}=h|_{\pi}$ is enriched with a directed graph structure $G_{\mathbf{h}}=(V,E)$, allowing the following implementation $\tilde{F}^{\theta}$ of $F^{\theta}$:
@@
The graph structure can be meticulously adjusted to implement various complexity-reducing techniques like Nyström approximation or integration domain truncation that effectively aim at a systematic reduction of the size of $\mathcal{N}_{\text{in}}(y)$; the naive implementation (12) is recovered for the case of a complete directed graph (with self-loops) for which $\mathcal{N}_{\text{in}}(y)=\pi$. Note that choosing $\mathcal{N}_{\text{in}}(y)$ as a strict subset of $\pi$ breaks the guaranteed smoothness in the GNO output.
@@ Appendix B Interpolation Graph Neural Operator
We detail our modifications of the GNO architecture. A generic Pytorch implementation is available at https://github.com/rwicl/operator-deep-smoothing-for-implied-volatility.
@@
In other words, we employ a Nyström approximation with nodes limited to the input data locations, which allows us to employ kernel functions with input skip connections. We set $G_{\mathbf{v}}=(\pi_{\text{out}},E)$, where
@@
  • $\mathcal{P}_{j}\colon\mathbb{R}^{c_{j}}\to\mathbb{R}^{\tilde{c}_{j}}$ and $\mathcal{Q}_{j}\colon\mathbb{R}^{\tilde{c}_{j+1}}\to\mathbb{R}^{c_{j+1}}$ are layer-individual lifting and projection, implemented as FNNs.
@@
    Both $\kappa^{W}_{j}$ and $\kappa^{b}_{j}$ are implemented as FNNs.
@@ C.1 Data
Our numerical experiments are based on the ”Option Quotes” dataset product available for purchase from the CBOE. Our version of the dataset contains relevant data features for S&P 500 Index options for the years 2012 through 2021 and is summarized on a 20-minute interval basis.
@@
We compute simple mids for options and underlying by averaging bid and ask quotes and use this aggregate as a reference price for all our subsequent computations. We calculate discount factors and forward prices from Put-Call parity using the industry standard technique based on linear regression. We compute time-to-expiry in units of one year as well as log-moneyness as defined in Section 2. We extract implied volatilities using the py-vollib-vectorized project available in the Python Package Index at the location https://pypi.org/project/py-vollib-vectorized/. py-vollib-vectorized implements a vectorized version of Jäckel, (2015)’s Let’s-be-rational state-of-the-art method for computing implied volatility. We discard all implied volatilities of in-the-money options, or, in other words, we compose our implied volatility surface from Put options for non-positive log-moneyness values and Call options for positive log-moneyness values.
@@ C.2 Model
We configure the modified GNO architecture introduced in Appendix B as follows:
  • We impose the following restriction on the in-neighborhood sets $\mathcal{N}_{\text{in}}(y)$:
    $\mathcal{N}_{\text{in}}(y)\subseteq\{(\rho_{l},z_{l})\in\pi\colon|\rho_{l}-\rho|\leq\overline{\rho}\}.$ (16)
    This choice stems from the consideration that implied volatility smoothing requires limited global informational exchange along the time-to-expiry axis. We then generate $\mathcal{N}_{\text{in}}(y)$ of certain size $K$ using a high-discrepancy subsampling heuristic detailed in the supplemented code material. Specifically, we use the values $\overline{\rho}=0.3$ and $K=50$.
@@
This configuration has been converged upon using manual experimentation and amounts to a total number of 102529 trainable parameters.
@@ C.3 Training
To ease notation we write $\hat{v}^{\theta}=\tilde{F}^{\theta}(\mathbf{v})$. We implement a Vega-weighted version of the fitting loss (8):
@@
$\mathcal{V}(\tau,x,v)=\partial_{v}\mathrm{BS}(\tau,k,v)=\varphi(d_{1}(\tau,k,v))\sqrt{\tau}.$ (17)
defined in Section 1 (adjusted for transformed coordinates).
For the implementation of the no-arbitrage penalization terms $\mathcal{L}_{\text{but}}$ and $\mathcal{L}_{\text{cal}}$, we first generate discretizations $\pi_{\rho}=\{\rho_{1},\dots,\rho_{m}\}$ and $\pi_{z}=\{z_{1},\dots,z_{n}\}$ of $[\rho_{\text{min}},\rho_{\text{max}}]$ and $[z_{\text{min}},z_{\text{max}}]$. We resolve the derivative terms $\partial_{z}v^{\theta}$ and $\partial_{z}^{2}v^{\theta}$ on the synthetic rectilinear grid $\pi=\pi_{\rho}\times\pi_{z}$ using (central) finite differences. Then, we translate $\mathcal{L}_{\text{but}}$ directly from (9) as
@@
Compared to a derivative based implementation, this implementation is independent of the scale and – in our empirical experiments – has provided an improved signal. Since the Nyström approximation employed by the graph neural operator (as well as the choice (16)) break the guaranteed smoothness of the operator output, we additionally introduce $\|\partial_{\rho}^{2}\hat{v}^{\theta}\|_{2}$ and $\|\partial^{2}_{z}\hat{v}^{\theta}\|_{2}$ as regularization terms:
@@
The particular weighting of the individual terms has been retrieved by manual experimentation, led by the findings of Ackerer et al., (2020).
@@
Table 2: Validation Loss
mean std min 25% 50% 75% max
$\mathcal{L}$ 0.062 0.032 0.029 0.053 0.059 0.064 0.506
$\mathcal{L}_{\text{fit}}$ 0.022 0.016 0.009 0.016 0.020 0.024 0.232
$\mathcal{L}_{\text{but}}$ 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\mathcal{L}_{\text{cal}}$ 0.000 0.001 0.000 0.000 0.000 0.000 0.008
$\mathcal{L}_{\text{reg-$\rho$}}$ 0.644 0.137 0.196 0.538 0.641 0.747 0.996
$\mathcal{L}_{\text{reg-$z$}}$ 0.143 0.254 0.048 0.083 0.101 0.134 5.154

### v3

@@ 5 Discussion
We provide a novel method for implied volatility smoothing, resulting from an application of our general operator approach for discretization-invariant data interpolation. The approach leverages a GNO to directly map given data – consistently across size and spatial arrangement – to smoothed surfaces, transcending classical parametric smoothing techniques. In the example of volatility smoothing, benefits include a massively simplified online calibration process.
@@
By moving the application of neural networks from the instance-by-instance level (Ackerer et al., 2020) to the “operator level”, we leverage the information contained in the entire training dataset for the smoothing of every single surface. In other words, our method “unlocks” large historical options datasets for volatility smoothing. We argue that our substantial outperformance against Ackerer et al. (2020) in the “Extrapolation-Test”-setting of the benchmark detailed in Table 5 of Appendix C.5 owes to this circumstance.
@@
Figure 3 makes the compression qualities of the operator approach apparent: We compute the entire historical timeseries using a single GNO instance, with around 100 thousand parameters. Evaluating the SVI benchmark, on the other hand, requires 61454 model instances (one per slice), or a total of 307270 parameters. A comparison with Ackerer et al. (2020), which for each smoothed surface trains a new neural network of around 5085 parameters, [^10] Computed as the sum of $120=3\times 40$ parameters for the input layer, three times $1640=41\times 40$ parameters for the hidden layers, $41$ parameters for the output layers, plus $4$ additional parameters of the SSVI prior and a scaling parameter. is striking: Smoothing of the CBOE dataset 2012–2021 at its 20-minute interval frequency would require more than 200 million parameters (more with rising frequency). At the same time, we expect our GNO to perform accurate smoothing over the entire training period and beyond (with regular finetuning), and our model instance remains fixed, even when moving to higher-frequency data.
@@
Compared to ad-hoc volatility parametrizations like SVI, our neural operator approach loses interpretability of parameters, which for some practitioners may be a stringent requirement. This disadvantage is generally shared by neural network based engineering solutions. Moreover, in some situations dimensionality reduction (even without interpretability of parameters) may be a desirable additional feature that is not directly achieved by our operator approach. Combining the VAE method (Bergeron et al., 2021) with our operator approach could lead to further promising potential applications of neural operators. Huang et al. (2024) introduces neural mappings, which generalize neural operators to mixed infinite-/finite-dimensionality for input or output spaces. This motivates a discretization-invariant GNO-based encoder, fit to handle raw incoming market data, and a classical decoder to extend the operator approach to a VAE-like architecture.
Reproducibility
Primarily, we ensure reproducibility by providing the codebase and model weights used to produce all results in this paper as part of the supplementary material. The codebase includes the data processing components, the GNO architecture, the loss functions, the error metrics, the production of the SVI benchmark, the notebooks used to train and evaluate the models (including hyperparameters and data splits), as well as the notebooks to produce the plots and tables in this paper. The full reproduction of results on intraday data (in particular of Figure 3) is contingent on access to the proprietary CBOE options data, which we are not allowed to provide. In fact, we have stripped the codebase from intermediate benchmarking artifacts that would expose the proprietary data (which some notebooks for the plots rely on). The dataset can be purchased from CBOE, but is expensive. The OptionMetrics end-of-day options data for the suite of indices considered in the final paragraph of Section 4, on the other hand, is more readily and freely available to researchers with subscriptions via the Wharton Research Data Services (WRDS) platform. The provided code allows to directly reproduce the experimental results, in particular, Table 1 and the plots in Appendix D.3. To do so, one would need to download the data from WRDS, persist it at prespecified location detailed in the codebase, and then run the respective notebooks, which automatically load the trained model weights.
To avoid any unclarities in our technique, the Pytorch implementation of our general graph neural operator architecture follows the mathematical definition given in Appendix B as closely as possible. Moreover, the concrete steps undertaken as part of our experiments are detailed in Appendix C:
  • Appendix C.1 gives a summary of the processing of the options data.
  • The hyperparameter configuration of our model finally employed in our experiments is detailed in Appendix C.2.
  • The loss functions and their weights are explicitly defined in Appendix C.3.
Ethics statement: data availability
The effectiveness of our method is fundamentally tied to the quality and frequency of the options data used for training. Our operator was trained using a proprietary dataset with 20-minute interval frequency, which may not be accessible to many researchers due to its cost. More easily available datasets, such as OptionMetrics data (e.g. freely available from the Wharton Research Data Services), usually only provide daily snapshots of volatility surfaces (end-of-day data). We acknowledge that training using such lower-frequency data may yield an operator that does not have the same smoothing performance as ours. This raises important ethical questions about data accessibility in financial engineering research, where high-quality, high-frequency data is often locked behind paywalls. In our case, the following strategies could potentially mitigate the resulting limitations: Augmenting low-frequency datasets with synthetically generated data based on established parametric models like SVI, or combining data from multiple indices to increase the effective sample size. Practitioners working with lower-frequency data might want to carefully evaluate these approaches and consider the relationship between their data sampling frequency and their intended use case.
@@ Acknowledgements
Access to data was made possible through the Clearify Project funding by an Imperial College SME engagement grant jointly with Zeliade Systems, as well as the MSc in Mathematics and Finance, Imperial College London. AJ gratefully acknowledges financial support from the EPSRC grants EP/W032643/1 and EP/T032146/1. RW was supported by the Department of Mathematics, Imperial College, through the Roth scholarship scheme. Computational resources and support were provided by the Imperial College Research Computing Service (Harvey, 2017). For the purpose of open access, the author(s) has applied a Creative Commons Attribution (CC BY) licence (where permitted by UKRI, ‘Open Government Licence’ or ‘Creative Commons Attribution No-derivatives (CC BY-ND) licence’ may be stated instead) to any Author Accepted Manuscript version arising.
@@ Appendix A Neural operators
We give a review of the kernel integral transform neural operator framework of Kovachki et al. (2023), and expand in more detail on its graph neural operator.
@@
Let $\mathcal{A}$ and $\mathcal{U}$ be input and output space of an operator learning problem, as introduced in Section 3.1. Then, $\mathcal{A}$ and $\mathcal{U}$ are Banach spaces of functions $D\to\mathbb{R}^{c_{\text{in}}}$ and $D\to\mathbb{R}^{c_{\text{out}}}$, respectively, where $D$ is a (bounded) domain in $\mathbb{R}^{d}$. The mathematical analysis of neural operators in Kovachki et al. (2023) summarized hereafter, as well as the definition of universality and discretization-invariance in Section 3.1, make use of the following terms and notations.
In the context of Kovachki et al. (2023), a domain is a bounded and connected open set that is topologically regular (in the sense that it is the interior of its closure). A domain $D$ is Lipschitz if its boundary locally is the graph of a Lipschitz continuous function defined on an open ball of $\mathbb{R}^{d-1}$. An open ball – for any metric space $\mathcal{X}=(\mathcal{X},d)$ – is the set
@@ A.1 Kernel integral neural operators
Theorem A.1 (Universal Approximation; Theorem 11 of Kovachki et al. (2023)).
@@
Kovachki et al. (2023) establishes the following.
Theorem A.2 (Discretization Invariance; Theorem 8 of Kovachki et al. (2023)).
Let $F^{\theta}\colon\mathcal{A}\to\mathcal{U}$ be a kernel integral neural operator, where $\mathcal{A}$ and $\mathcal{U}$ both continuously embed into $C(\overline{D})$. Then, the implementation of $F^{\theta}$ based on equation 12 is discretization-invariant as defined in Section 3.1.
equation 12 suggests the straightforward (quasi) Monte-Carlo inspired implementation
@@ A.2 Graph neural operators
The curse of dimensionality makes the direct implementation equation 13 prohibitively expensive in practice. Instead, Anandkumar et al. (2020) introduces graph neural operators (or, GNOs, for short) which replace the kernel integral operation at the heart of the framework by a sum approximation and organizes the constituent terms using a directed graph structure: The discretization $\pi=\{x_{1},\dots,x_{m}\}$ of the input data $\mathbf{h}=h|_{\pi}$ is enriched with a directed graph structure $G_{\mathbf{h}}=(V,E)$, allowing the following implementation $\tilde{F}^{\theta}$ of $F^{\theta}$:
@@
The graph structure can be meticulously adjusted to implement various complexity-reducing techniques like Nyström approximation or integration domain truncation that effectively aim at a systematic reduction of the size of $\mathcal{N}_{\text{in}}(y)$; the naive implementation equation 12 is recovered for the case of a complete directed graph (with self-loops) for which $\mathcal{N}_{\text{in}}(y)=\pi$. Note that choosing $\mathcal{N}_{\text{in}}(y)$ as a strict subset of $\pi$ breaks the guaranteed smoothness in the GNO output.
@@ Appendix B Interpolation graph neural operator
We detail our modifications of the GNO architecture.
@@
In other words, we employ a Nyström approximation with nodes limited to the input data locations. This is an important prerequisite to enable the use of the GNO architecture for interpolation tasks (or, more generally phrased, allows us to employ kernel functions with input skip connections). We set $G_{\mathbf{v}}=(\pi_{\text{out}},E)$, where
@@
  • $\mathcal{P}_{j}\colon\mathbb{R}^{c_{j}}\to\mathbb{R}^{\tilde{c}_{j}}$ and $\mathcal{Q}_{j}\colon\mathbb{R}^{\tilde{c}_{j+1}}\to\mathbb{R}^{c_{j+1}}$ are layer-individual lifting and projection, in view of A.1 implemented simply as FNNs.
@@
    Both $\kappa^{W}_{j}$ and $\kappa^{b}_{j}$ are implemented as FNNs in our case, again to satisfy the requirements of A.1 and to keep things simple.
@@ C.1 Data
Our numerical experiments are based on the “Option Quotes” dataset product available for purchase from the CBOE. Our version of the dataset contains relevant data features for S&P 500 Index options for the years 2012 through 2021 and is summarized on a 20-minute interval basis.
@@
We compute simple mids for options and underlying by averaging bid and ask quotes and use this aggregate as a reference price for all our subsequent computations. We calculate discount factors and forward prices from Put-Call parity using the industry standard technique based on linear regression. We compute time-to-expiry in units of one year as well as log-moneyness as defined in Section 2. We extract implied volatilities using the py-vollib-vectorized project available in the Python Package Index at the location https://pypi.org/project/py-vollib-vectorized/. py-vollib-vectorized implements a vectorized version of Jäckel (2015)’s Let’s-be-rational state-of-the-art method for computing implied volatility. We discard all implied volatilities of in-the-money options, or, in other words, we compose our implied volatility surface from Put options for non-positive log-moneyness values and Call options for positive log-moneyness values.
@@ C.2 Model
We proceed to detail the hyperparameter configuration of the modified GNO architecture introduced in Appendix B.
The choice of in-neighborhoods
The construction of the in-neighborhood sets for the graph neural operator is a crucial hyperparameter choice, fundamentally dictating the computation routes (and thus complexity) of the forward pass of the model. We already explained in Appendix B that we employ a Nyström approximation with subsampling from the input data nodes, to unlock the GNO for interpolation tasks. Additionally, we employ truncation. Truncation limits the spatial extent of the in-neighborhoods and is a way to incorporate information about the locality structure of the learning task at hand directly into the graph neural operator architecture. Since implied volatility smoothing requires limited global informational exchange along the time-to-expiry axis, we impose the following restriction on the in-neighborhood sets $\mathcal{N}_{\text{in}}(y)$:
$\mathcal{N}_{\text{in}}(y)\subseteq\overline{\mathcal{N}_{\text{in}}}(y),$ (16)
where
$\overline{\mathcal{N}_{\text{in}}}(\rho_{y},z_{y})=\{(\rho_{l},z_{l})\in\pi\colon|\rho_{l}-\rho_{y}|\leq\overline{\rho}\}$ (17)
is the set of all available options $(\rho_{l},z_{l})$ contained in the slices with a time-to-expiry $\rho_{l}$ close than $\bar{\rho}$ to the time-to-expiry $\rho_{y}$ of $y$. We explain our reasoning more precisely:
  • The input data for volatility smoothing is not arbitrarily scattered over the $(\rho,z)$-domain, but arranged as dense $z$-slices that are sparseley distributed along the $\rho$-axis (compare Figure 5 as well as three more examples pictured in Figure 1(b)). Condition 2.1(i) of Theorem 2.1 imposes monotonicity of the output surface along the time-to-expiry axis. This constraint is inherently “local”: To generate a compliant output surface, it is sufficient for the hidden states at a given output location to receive information from their immediate neighboring slices. [^11] A collection of slices that is monotonously increasing in pairs is montonously increasing as a whole. We computed the maximum distance (with respect to $\rho$-coordinates) between slices over our entire dataset as $\Delta_{\text{max}}\rho\approx 0.269$, which is thus established as a lower bound for $\bar{\rho}$, and finally explains our choice $\bar{\rho}=0.3$. We note that – because we use three hidden GNO layers (see below) – the domain of influence of each input point ultimately is unrestricted: The compositional structure allows information to travel slice to slice in steps of length $\bar{\rho}=0.3$, which amounts to a total distance of $4\times 0.3=1.2$. This exceeds the size of the considered domain $D=(\rho_{\text{min}},\rho_{\text{max}})\times(z_{\text{min}},z_{\text{max}})$ in the direction of the $\rho$-axis. It is therefore not motivated to increase the level of $\bar{\rho}$.
  • We do not perform a similar truncation in the direction of the log-moneyness axis. In particular, this allows all points in any given slice to connect with each other indiscriminately, which in view of the nonlinear shape constraint 2.1(ii) of Theorem 2.1 is motivated. Moreover, such truncation would limit extrapolation distance in $z$-direction.
To concretely compute $\mathcal{N}_{\text{in}}(y)$ for given $y$ on the basis of equation 16, we employ the following low-discrepancy subsampling heuristic, parametrized by the hyperparameter $K$: First, we compute $\overline{\mathcal{N}_{\text{in}}}(y)$ and convert it to a sequence by sorting it by two-dimensional Euclidean distance to $y$ in ascending order. Of this sequence we take every $k$-th element, where $k$ is the largest step size such that the final number of nodes $\mathcal{N}_{\text{in}}(y)$ does not exceed $K$. This gives us $\mathcal{N}_{\text{in}}(y)$. Note that, by sorting $\overline{\mathcal{N}_{\text{in}}}(y)$ and performing a “sparse” selection, we promote low-discrepancy properties for $\mathcal{N}_{\text{in}}(y)$, which intuitively aid the convergence properties of the kernel integrals (compare 5).
The hyperparameter $K$, finally, constitutes an upper bound on the size of the $\mathcal{N}_{\text{in}}(y)$. It allows us to control the computational complexity of the model, in a trade-off, of course, with the expressivity of the GNO. After manual experimentation, we settle on a value of $K=50$ and perform an ablation study in Appendix C.6 to validate our choice.
Figure 5: Graph construction: Given data $\mathbf{v}=(v_{x})_{x\in\pi}$ (S&P 500 on 10.05.2012 at 16:10:00), generate in-neighborhood for output point $y$ by subsampling from $\overline{\mathcal{N}_{\text{in}}}(\rho_{y},z_{y})$ by first sorting and then taking a sparse selection, promoting low-discrepancy.
GNO layers and kernels
The below choices amount to a total number of 102529 trainable parameters.
@@ C.3 Training
To ease notation we write $\hat{v}^{\theta}=\tilde{F}^{\theta}(\mathbf{v})$. We implement a Vega-weighted version of the fitting loss equation 8:
@@
$\mathcal{V}(\rho,z,v)=\partial_{v}\mathrm{BS}(\tau,k,v)=\varphi(d_{1}(\tau,k,v))\sqrt{\tau}.$ (18)
For the implementation of the no-arbitrage penalization terms $\mathcal{L}_{\text{but}}$ and $\mathcal{L}_{\text{cal}}$, we first generate discretizations $\pi_{\rho}=\{\rho_{1},\dots,\rho_{m}\}$ and $\pi_{z}=\{z_{1},\dots,z_{n}\}$ of $[\rho_{\text{min}},\rho_{\text{max}}]$ and $[z_{\text{min}},z_{\text{max}}]$. We resolve the derivative terms $\partial_{z}v^{\theta}$ and $\partial_{z}^{2}v^{\theta}$ on the synthetic rectilinear grid $\pi=\pi_{\rho}\times\pi_{z}$ using (central) finite differences. Then, we translate $\mathcal{L}_{\text{but}}$ directly from equation 9 as
@@
Compared to a derivative based implementation, this implementation is independent of the scale and – in our empirical experiments – has provided an improved signal. Since the Nyström approximation employed by the graph neural operator (as well as the choice equation 16) break the guaranteed smoothness of the operator output, we additionally introduce $\|\partial_{\rho}^{2}\hat{v}^{\theta}\|_{2}$ and $\|\partial^{2}_{z}\hat{v}^{\theta}\|_{2}$ as regularization terms:
@@
The particular weighting of the individual terms has initially been retrieved by manual experimentation, led by the findings of Ackerer et al. (2020). To additionally validate our choices, we perform an ablation study in Appendix C.7.
@@
Table 2: Validation loss.
mean std 1% 25% 50% 75% 99%
$\mathcal{L}$ 0.0591 0.0807 0.0351 0.0450 0.0506 0.0578 0.1489
$\mathcal{L}_{\text{fit}}$ 0.0182 0.0182 0.0066 0.0121 0.0162 0.0203 0.0479
$\mathcal{L}_{\text{but}}$ 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
$\mathcal{L}_{\text{cal}}$ 0.0001 0.0006 0.0000 0.0000 0.0000 0.0000 0.0001
$\mathcal{L}_{\text{reg-$r$}}$ 0.8567 2.4065 0.3280 0.4702 0.5788 0.7703 2.7578
$\mathcal{L}_{\text{reg-$z$}}$ 0.7610 0.1403 0.4815 0.6488 0.7590 0.8679 1.0812